OpenStudy (anonymous):

Let f(x) = 5x + 12. Find f^–1(x). Help I don't understand how to do this:(

OpenStudy (anonymous):

suppose you wanted to solve \[5x+12=10\] what would you do?

OpenStudy (anonymous):

\[5x+12=10\] 1) subtract 12 \[5x=-2\] 2) divide by 5 \[x=-\frac{2}{5}\] so in other words if you know the output is 10, you know x must have been \[-\frac{2}{5}\] this is how the inverse works. it tells you what the input is if you know the output

OpenStudy (anonymous):

so the inverse says 1) subtract 12 2) divide by 5 and we can write \[f^{-1}(x)=\frac{x-12}{5}\]

OpenStudy (anonymous):

another (identical) way to think about it is this. \[f(x)=5x+12\] says 1) multiply by 5, 2) add 12 inverse will do the reverse operations in reverse order, so again it is 1) subtract 12 2) divide by 5

OpenStudy (anonymous):

okay, thank you:)

OpenStudy (anonymous):

and finally (they way you might have seen) is take \[y=5x+12\] switch x and y to write \[x=5y+12\] solve for y and get \[x-12=5y\] \[\frac{x-12}{5}+y\] and so \[f^{-1}(x)=y=\frac{x-12}{5}\]

OpenStudy (anonymous):

beat that one to death, huh?

OpenStudy (anonymous):

haha yes, thank you! I needed it:))